# DiffSampling: Enhancing Diversity and Accuracy in Neural Text Generation

### v4

@@ D.4 Ablation Study on the Dynamic Upper Bound
Top-$p$ $0.24_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.01}$ $0.59_{\pm.01}$ $0.33_{\pm.00}$ $0.66_{\pm.01}$ $0.30_{\pm.01}$ $0.16_{\pm.00}$ $0.35_{\pm.01}$ $0.50_{\pm.01}$ $0.73_{\pm.01}$ $0.66_{\pm.01}$ $0.76_{\pm.01}$ $0.55_{\pm.01}$
$\eta$-sampling $0.23_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.00}$ $0.64_{\pm.01}$ $0.35_{\pm.00}$ $0.70_{\pm.01}$ $0.31_{\pm.01}$ $0.15_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.76_{\pm.01}$ $0.68_{\pm.00}$ $0.80_{\pm.01}$ $0.56_{\pm.01}$
Locally typical $0.24_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.01}$ $0.59_{\pm.01}$ $0.33_{\pm.00}$ $0.66_{\pm.01}$ $0.30_{\pm.01}$ $0.16_{\pm.00}$ $0.35_{\pm.01}$ $0.50_{\pm.01}$ $0.73_{\pm.01}$ $0.66_{\pm.01}$ $0.76_{\pm.01}$ $0.55_{\pm.01}$
Min-$p$ $0.25_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.59_{\pm.01}$ $0.30_{\pm.00}$ $0.65_{\pm.01}$ $0.28_{\pm.01}$ $0.19_{\pm.00}$ $0.45_{\pm.01}$ $0.65_{\pm.01}$ $0.63_{\pm.01}$ $0.47_{\pm.01}$ $0.61_{\pm.01}$ $0.39_{\pm.01}$
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.22_{\pm.00}$ $0.47_{\pm.01}$ $0.21_{\pm.01}$ $0.20_{\pm.00}$ $0.51_{\pm.00}$ $0.77_{\pm.01}$ $0.40_{\pm.01}$ $0.21_{\pm.00}$ $0.33_{\pm.01}$ $0.19_{\pm.01}$
DiffS.-lb $0.24_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.00}$ $0.60_{\pm.01}$ $0.33_{\pm.00}$ $0.67_{\pm.01}$ $0.30_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.49_{\pm.01}$ $0.73_{\pm.01}$ $0.67_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
DiffS.-minp $0.25_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.59_{\pm.01}$ $0.30_{\pm.00}$ $0.65_{\pm.01}$ $0.28_{\pm.01}$ $0.19_{\pm.00}$ $0.45_{\pm.01}$ $0.64_{\pm.01}$ $0.63_{\pm.01}$ $0.48_{\pm.01}$ $0.61_{\pm.01}$ $0.39_{\pm.01}$
@@
Top-$p$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.54_{\pm.01}$ $0.26_{\pm.00}$ $0.57_{\pm.01}$ $0.24_{\pm.01}$ $0.19_{\pm.00}$ $0.47_{\pm.01}$ $0.71_{\pm.01}$ $0.52_{\pm.01}$ $0.38_{\pm.01}$ $0.49_{\pm.01}$ $0.31_{\pm.01}$
$\eta$-sampling $0.25_{\pm.00}$ $0.53_{\pm.01}$ $0.69_{\pm.00}$ $0.56_{\pm.01}$ $0.28_{\pm.00}$ $0.60_{\pm.01}$ $0.25_{\pm.01}$ $0.19_{\pm.00}$ $0.46_{\pm.01}$ $0.68_{\pm.01}$ $0.56_{\pm.01}$ $0.43_{\pm.01}$ $0.53_{\pm.01}$ $0.34_{\pm.01}$
Locally typical $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.54_{\pm.01}$ $0.26_{\pm.00}$ $0.57_{\pm.01}$ $0.24_{\pm.01}$ $0.19_{\pm.00}$ $0.47_{\pm.01}$ $0.71_{\pm.01}$ $0.52_{\pm.01}$ $0.38_{\pm.01}$ $0.49_{\pm.01}$ $0.31_{\pm.01}$
Min-$p$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.53_{\pm.01}$ $0.26_{\pm.00}$ $0.56_{\pm.01}$ $0.24_{\pm.01}$ $0.19_{\pm.00}$ $0.48_{\pm.01}$ $0.73_{\pm.01}$ $0.50_{\pm.01}$ $0.35_{\pm.01}$ $0.46_{\pm.01}$ $0.28_{\pm.01}$
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.21_{\pm.00}$ $0.46_{\pm.01}$ $0.19_{\pm.01}$ $0.20_{\pm.00}$ $0.51_{\pm.00}$ $0.77_{\pm.01}$ $0.39_{\pm.01}$ $0.20_{\pm.00}$ $0.31_{\pm.01}$ $0.18_{\pm.01}$
DiffS.-lb $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.55_{\pm.01}$ $0.28_{\pm.00}$ $0.60_{\pm.01}$ $0.26_{\pm.01}$ $0.19_{\pm.00}$ $0.46_{\pm.01}$ $0.68_{\pm.01}$ $0.56_{\pm.01}$ $0.43_{\pm.01}$ $0.53_{\pm.01}$ $0.34_{\pm.01}$
DiffS.-minp $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.55_{\pm.01}$ $0.27_{\pm.00}$ $0.59_{\pm.01}$ $0.25_{\pm.01}$ $0.19_{\pm.00}$ $0.47_{\pm.01}$ $0.70_{\pm.01}$ $0.53_{\pm.01}$ $0.39_{\pm.01}$ $0.50_{\pm.01}$ $0.32_{\pm.01}$
@@
Top-$p$ $0.06_{\pm.00}$ $0.16_{\pm.01}$ $0.30_{\pm.01}$ $0.65_{\pm.01}$ $0.62_{\pm.01}$ $0.90_{\pm.01}$ $76_{\pm.01} 0.0$$3_{\pm.00} 0.0$$07_{\pm.00}$ $0.20_{\pm.00}$ $0.80_{\pm.01}$ $0.71_{\pm.00}$ $0.97_{\pm.00}$ $0.90_{\pm.00}$
$\eta$-sampling $0.09_{\pm.00}$ $0.25_{\pm.01}$ $0.38_{\pm.01}$ $0.71_{\pm.01}$ $0.61_{\pm.01}$ $0.88_{\pm.01}$ $65_{\pm.01} 0.0$$3_{\pm.00} 0.0$$07_{\pm.00}$ $0.20_{\pm.00}$ $0.81_{\pm.01}$ $0.72_{\pm.00}$ $0.97_{\pm.00}$ $0.89_{\pm.00}$
Locally typical $0.06_{\pm.00}$ $0.15_{\pm.01}$ $0.29_{\pm.01}$ $0.68_{\pm.01}$ $0.59_{\pm.01}$ $0.90_{\pm.01}$ $77_{\pm.01} 0.0$$2_{\pm.00} 0.0$$06_{\pm.00}$ $0.19_{\pm.00}$ $0.83_{\pm.01}$ $0.65_{\pm.00}$ $0.97_{\pm.00}$ $0.91_{\pm.00}$
Min-$p$ $0.23_{\pm.00}$ $0.52_{\pm.01}$ $0.66_{\pm.00}$ $0.63_{\pm.01}$ $0.36_{\pm.00}$ $0.72_{\pm.01}$ $0.32_{\pm.01}$ $0.17_{\pm.00}$ $0.38_{\pm.01}$ $0.53_{\pm.01}$ $0.77_{\pm.01}$ $0.61_{\pm.00}$ $0.78_{\pm.01}$ $0.51_{\pm.01}$
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.21_{\pm.01}$ $0.20_{\pm.00}$ $0.51_{\pm.00}$ $0.77_{\pm.01}$ $0.40_{\pm.01}$ $0.21_{\pm.00}$ $0.34_{\pm.01}$ $0.19_{\pm.01}$
DiffS.-lb $0.23_{\pm.00}$ $0.51_{\pm.01}$ $0.66_{\pm.01}$ $0.64_{\pm.01}$ $0.37_{\pm.00}$ $0.72_{\pm.01}$ $0.33_{\pm.01}$ $0.11_{\pm.00}$ $0.23_{\pm.01}$ $0.34_{\pm.01}$ $0.77_{\pm.01}$ $0.79_{\pm.00}$ $0.89_{\pm.00}$ $0.70_{\pm.01}$
DiffS.-minp $0.24_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.60_{\pm.01}$ $0.33_{\pm.00}$ $0.68_{\pm.01}$ $0.30_{\pm.01}$ $0.18_{\pm.00}$ $0.43_{\pm.01}$ $0.61_{\pm.01}$ $0.69_{\pm.01}$ $0.52_{\pm.00}$ $0.68_{\pm.01}$ $0.44_{\pm.01}$
@@
Top-$p$ $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.14_{\pm.00}$ $0.65_{\pm.01}$ $0.61_{\pm.00}$ $0.94_{\pm.00}$ $94_{\pm.00} 0.0$$1_{\pm.00} 0.0$$03_{\pm.00}$ $0.16_{\pm.00}$ $0.82_{\pm.01}$ $0.57_{\pm.00}$ $0.98_{\pm.00}$ $0.94_{\pm.00}$
$\eta$-sampling $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.14_{\pm.00}$ $0.65_{\pm.01}$ $0.63_{\pm.00}$ $0.94_{\pm.00}$ $94_{\pm.00} 0.0$$1_{\pm.00} 0.0$$03_{\pm.00}$ $0.15_{\pm.00}$ $0.81_{\pm.01}$ $0.59_{\pm.00}$ $0.98_{\pm.00}$ $0.94_{\pm.00}$
Locally typical $00_{\pm.00} 0.0$$03_{\pm.00}$ $0.14_{\pm.00}$ $0.73_{\pm.01}$ $0.58_{\pm.00}$ $0.95_{\pm.00}$ $95_{\pm.00} 0.0$$1_{\pm.00} 0.0$$03_{\pm.00}$ $0.15_{\pm.00}$ $0.82_{\pm.01}$ $0.55_{\pm.00}$ $0.98_{\pm.00}$ $0.94_{\pm.00}$
Min-$p$ $0.21_{\pm.00}$ $0.49_{\pm.01}$ $0.64_{\pm.01}$ $0.64_{\pm.01}$ $0.42_{\pm.00}$ $0.77_{\pm.01}$ $0.37_{\pm.01}$ $0.13_{\pm.00}$ $0.30_{\pm.01}$ $0.42_{\pm.01}$ $0.81_{\pm.01}$ $0.72_{\pm.00}$ $0.89_{\pm.00}$ $0.63_{\pm.01}$
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.22_{\pm.00}$ $0.49_{\pm.01}$ $0.21_{\pm.01}$ $0.20_{\pm.00}$ $0.50_{\pm.00}$ $0.77_{\pm.01}$ $0.41_{\pm.01}$ $0.21_{\pm.00}$ $0.34_{\pm.01}$ $0.20_{\pm.01}$
DiffS.-lb $0.22_{\pm.00}$ $0.50_{\pm.01}$ $0.64_{\pm.01}$ $0.61_{\pm.01}$ $0.41_{\pm.00}$ $0.74_{\pm.01}$ $35_{\pm.01} 0.0$$08_{\pm.00}$ $0.19_{\pm.01}$ $0.29_{\pm.01}$ $0.78_{\pm.01}$ $0.82_{\pm.00}$ $0.92_{\pm.00}$ $0.76_{\pm.01}$
DiffS.-minp $0.24_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.63_{\pm.01}$ $0.33_{\pm.00}$ $0.70_{\pm.01}$ $0.30_{\pm.01}$ $0.18_{\pm.00}$ $0.42_{\pm.01}$ $0.60_{\pm.01}$ $0.70_{\pm.01}$ $0.54_{\pm.00}$ $0.71_{\pm.01}$ $0.46_{\pm.01}$
@@
Top-$p$ $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.76_{\pm.01}$ $0.60_{\pm.00}$ $0.96_{\pm.00}$ $95_{\pm.00} 0.0$$0_{\pm.00} 0.0$$03_{\pm.00}$ $0.13_{\pm.00}$ $0.79_{\pm.01}$ $0.57_{\pm.00}$ $0.98_{\pm.00}$ $0.95_{\pm.00}$
$\eta$-sampling $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.61_{\pm.00}$ $0.96_{\pm.00}$ $95_{\pm.00} 0.0$$0_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.76_{\pm.01}$ $0.60_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$
Locally typical $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.76_{\pm.01}$ $0.60_{\pm.00}$ $0.96_{\pm.00}$ $95_{\pm.00} 0.0$$0_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.76_{\pm.01}$ $0.58_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$
Min-$p$ $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.61_{\pm.00}$ $0.96_{\pm.00}$ $95_{\pm.00} 0.0$$0_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.76_{\pm.01}$ $0.60_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.45_{\pm.01}$ $0.22_{\pm.00}$ $0.50_{\pm.01}$ $0.21_{\pm.01}$ $0.20_{\pm.00}$ $0.50_{\pm.00}$ $0.77_{\pm.01}$ $0.41_{\pm.01}$ $0.21_{\pm.00}$ $0.35_{\pm.01}$ $0.20_{\pm.01}$
DiffS.-lb $0.20_{\pm.00}$ $0.46_{\pm.01}$ $0.61_{\pm.01}$ $0.65_{\pm.01}$ $0.44_{\pm.01}$ $0.79_{\pm.01}$ $40_{\pm.01} 0.0$$04_{\pm.00}$ $0.12_{\pm.00}$ $0.24_{\pm.01}$ $0.79_{\pm.01}$ $0.80_{\pm.00}$ $0.95_{\pm.00}$ $0.83_{\pm.00}$
DiffS.-minp $0.23_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.00}$ $0.62_{\pm.01}$ $0.35_{\pm.00}$ $0.72_{\pm.01}$ $0.32_{\pm.01}$ $0.17_{\pm.00}$ $0.40_{\pm.01}$ $0.56_{\pm.01}$ $0.75_{\pm.01}$ $0.58_{\pm.00}$ $0.77_{\pm.01}$ $0.50_{\pm.01}$

### v5

@@ F.2 Extreme Summarization
Top-$p$ $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.59_{\pm.01}$ $0.36_{\pm.01}$ $0.47_{\pm.01}$ $0.66_{\pm.01}$ $0.41_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.72_{\pm.01}$ $0.66_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
$\eta$-sampling $0.20_{\pm.00}$ $0.45_{\pm.01}$ $0.58_{\pm.01}$ $0.38_{\pm.01}$ $0.49_{\pm.01}$ $0.69_{\pm.01}$ $0.43_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.75_{\pm.01}$ $0.67_{\pm.00}$ $0.80_{\pm.01}$ $0.56_{\pm.01}$
Locally typical $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.59_{\pm.01}$ $0.36_{\pm.01}$ $0.47_{\pm.01}$ $0.66_{\pm.01}$ $0.41_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.72_{\pm.01}$ $0.66_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
Min-$p$ $0.22_{\pm.00}$ $0.46_{\pm.01}$ $0.61_{\pm.01}$ $0.36_{\pm.01}$ $0.43_{\pm.01}$ $0.64_{\pm.01}$ $0.38_{\pm.01}$ $0.20_{\pm.00}$ $0.44_{\pm.01}$ $0.63_{\pm.01}$ $0.65_{\pm.01}$ $0.47_{\pm.01}$ $0.62_{\pm.01}$ $0.39_{\pm.01}$
DiffS.-cut $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.45_{\pm.01}$ $0.23_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.00}$ $0.38_{\pm.01}$ $0.19_{\pm.00}$ $0.32_{\pm.01}$ $0.17_{\pm.01}$
DiffS.-lb $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.59_{\pm.01}$ $0.37_{\pm.01}$ $0.47_{\pm.01}$ $0.67_{\pm.01}$ $0.41_{\pm.01}$ $0.16_{\pm.00}$ $0.34_{\pm.01}$ $0.48_{\pm.01}$ $0.72_{\pm.01}$ $0.66_{\pm.01}$ $0.77_{\pm.01}$ $0.55_{\pm.01}$
DiffS.-minp $0.22_{\pm.00}$ $0.46_{\pm.01}$ $0.60_{\pm.01}$ $0.35_{\pm.01}$ $0.43_{\pm.01}$ $0.64_{\pm.01}$ $0.38_{\pm.01}$ $0.20_{\pm.00}$ $0.44_{\pm.01}$ $0.62_{\pm.01}$ $0.65_{\pm.01}$ $0.47_{\pm.01}$ $0.63_{\pm.01}$ $0.39_{\pm.01}$
@@
Top-$p$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.62_{\pm.01}$ $0.38_{\pm.01}$ $0.34_{\pm.01}$ $0.56_{\pm.01}$ $0.30_{\pm.01}$ $0.20_{\pm.00}$ $0.47_{\pm.01}$ $0.68_{\pm.01}$ $0.53_{\pm.01}$ $0.37_{\pm.01}$ $0.50_{\pm.01}$ $0.30_{\pm.01}$
$\eta$-sampling $0.22_{\pm.00}$ $0.47_{\pm.01}$ $0.62_{\pm.01}$ $0.38_{\pm.01}$ $0.38_{\pm.01}$ $0.59_{\pm.01}$ $0.33_{\pm.01}$ $0.20_{\pm.00}$ $0.46_{\pm.01}$ $0.66_{\pm.01}$ $0.59_{\pm.01}$ $0.42_{\pm.01}$ $0.55_{\pm.01}$ $0.33_{\pm.01}$
Locally typical $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.62_{\pm.01}$ $0.38_{\pm.01}$ $0.34_{\pm.01}$ $0.56_{\pm.01}$ $0.30_{\pm.01}$ $0.20_{\pm.00}$ $0.47_{\pm.01}$ $0.68_{\pm.01}$ $0.53_{\pm.01}$ $0.37_{\pm.01}$ $0.50_{\pm.01}$ $0.30_{\pm.01}$
Min-$p$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.39_{\pm.01}$ $0.34_{\pm.01}$ $0.55_{\pm.01}$ $0.30_{\pm.01}$ $0.21_{\pm.00}$ $0.47_{\pm.01}$ $0.69_{\pm.01}$ $0.50_{\pm.01}$ $0.33_{\pm.01}$ $0.46_{\pm.01}$ $0.28_{\pm.01}$
DiffS.-cut $0.23_{\pm.00}$ $0.49_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.24_{\pm.01}$ $0.43_{\pm.01}$ $0.22_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.00}$ $0.37_{\pm.01}$ $0.18_{\pm.00}$ $0.30_{\pm.01}$ $0.16_{\pm.01}$
DiffS.-lb $0.22_{\pm.00}$ $0.47_{\pm.01}$ $0.62_{\pm.01}$ $0.39_{\pm.01}$ $0.38_{\pm.01}$ $0.59_{\pm.01}$ $0.32_{\pm.01}$ $0.20_{\pm.00}$ $0.45_{\pm.01}$ $0.65_{\pm.01}$ $0.59_{\pm.01}$ $0.42_{\pm.01}$ $0.55_{\pm.01}$ $0.34_{\pm.01}$
DiffS.-minp $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.62_{\pm.01}$ $0.39_{\pm.01}$ $0.37_{\pm.01}$ $0.58_{\pm.01}$ $0.32_{\pm.01}$ $0.20_{\pm.00}$ $0.47_{\pm.01}$ $0.67_{\pm.01}$ $0.55_{\pm.01}$ $0.38_{\pm.01}$ $0.52_{\pm.01}$ $0.31_{\pm.01}$
@@
Top-$p$ $0.07_{\pm.00}$ $0.18_{\pm.01}$ $0.30_{\pm.01}$ $0.64_{\pm.01}$ $0.72_{\pm.00}$ $0.88_{\pm.01}$ $77_{\pm.01} 0.0$$3_{\pm.00} 0.0$$08_{\pm.00}$ $0.19_{\pm.00}$ $9_{\pm.01} 0.77±.0$$7_{\pm.00}$ $0.97_{\pm.00}$ $0.88_{\pm.00}$
$\eta$-sampling $0.10_{\pm.00}$ $0.25_{\pm.01}$ $0.38_{\pm.01}$ $0.63_{\pm.01}$ $0.70_{\pm.00}$ $0.86_{\pm.01}$ $69_{\pm.01} 0.0$$3_{\pm.00} 0.0$$08_{\pm.00}$ $0.19_{\pm.00}$ $9_{\pm.01} 0.77±.0$$7_{\pm.00}$ $0.96_{\pm.00}$ $0.88_{\pm.00}$
Locally typical $0.06_{\pm.00}$ $0.17_{\pm.01}$ $0.30_{\pm.01}$ $0.69_{\pm.01}$ $0.70_{\pm.00}$ $0.89_{\pm.01}$ $77_{\pm.01} 0.0$$2_{\pm.00} 0.0$$07_{\pm.00}$ $0.19_{\pm.00}$ $0.82_{\pm.01}$ $0.72_{\pm.00}$ $0.97_{\pm.00}$ $0.90_{\pm.00}$
Min-$p$ $0.20_{\pm.00}$ $0.44_{\pm.01}$ $0.58_{\pm.01}$ $0.37_{\pm.01}$ $0.50_{\pm.01}$ $0.71_{\pm.01}$ $0.45_{\pm.01}$ $0.17_{\pm.00}$ $0.38_{\pm.01}$ $0.53_{\pm.01}$ $0.78_{\pm.01}$ $0.60_{\pm.00}$ $0.79_{\pm.01}$ $0.51_{\pm.01}$
DiffS.-cut $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.46_{\pm.01}$ $0.23_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.01}$ $0.38_{\pm.01}$ $0.19_{\pm.00}$ $0.33_{\pm.01}$ $0.18_{\pm.01}$
DiffS.-lb $0.20_{\pm.00}$ $0.43_{\pm.01}$ $0.57_{\pm.01}$ $0.40_{\pm.01}$ $0.52_{\pm.01}$ $0.71_{\pm.01}$ $0.45_{\pm.01}$ $0.11_{\pm.00}$ $0.23_{\pm.01}$ $0.34_{\pm.01}$ $0.75_{\pm.01}$ $0.79_{\pm.00}$ $0.89_{\pm.00}$ $0.70_{\pm.01}$
DiffS.-minp $0.21_{\pm.00}$ $0.46_{\pm.01}$ $0.60_{\pm.01}$ $0.35_{\pm.01}$ $0.45_{\pm.01}$ $0.67_{\pm.01}$ $0.41_{\pm.01}$ $0.19_{\pm.00}$ $0.43_{\pm.01}$ $0.60_{\pm.01}$ $0.70_{\pm.01}$ $0.52_{\pm.00}$ $0.70_{\pm.01}$ $0.44_{\pm.01}$
@@
Top-$p$ $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.14_{\pm.00}$ $0.63_{\pm.01}$ $0.65_{\pm.00}$ $0.92_{\pm.01}$ $91_{\pm.00} 0.0$$1_{\pm.00} 0.0$$03_{\pm.00}$ $0.15_{\pm.00}$ $0.81_{\pm.01}$ $0.65_{\pm.00}$ $0.98_{\pm.00}$ $0.94_{\pm.00}$
$\eta$-sampling $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.13_{\pm.00}$ $0.65_{\pm.01}$ $0.67_{\pm.00}$ $0.92_{\pm.01}$ $92_{\pm.00} 0.0$$1_{\pm.00} 0.0$$03_{\pm.00}$ $0.14_{\pm.00}$ $0.80_{\pm.01}$ $0.66_{\pm.00}$ $0.98_{\pm.00}$ $0.93_{\pm.00}$
Locally typical $00_{\pm.00} 0.0$$03_{\pm.00}$ $0.14_{\pm.00}$ $0.71_{\pm.01}$ $0.63_{\pm.00}$ $0.93_{\pm.00}$ $92_{\pm.00} 0.0$$1_{\pm.00} 0.0$$03_{\pm.00}$ $0.15_{\pm.00}$ $0.83_{\pm.01}$ $0.63_{\pm.00}$ $0.98_{\pm.00}$ $0.94_{\pm.00}$
Min-$p$ $0.19_{\pm.00}$ $0.42_{\pm.01}$ $0.55_{\pm.01}$ $0.41_{\pm.01}$ $0.56_{\pm.01}$ $0.76_{\pm.01}$ $0.50_{\pm.01}$ $0.13_{\pm.00}$ $0.30_{\pm.01}$ $0.43_{\pm.01}$ $0.82_{\pm.01}$ $0.72_{\pm.00}$ $0.89_{\pm.00}$ $0.62_{\pm.01}$
DiffS.-cut $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.46_{\pm.01}$ $0.24_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.01}$ $0.38_{\pm.01}$ $0.19_{\pm.00}$ $0.33_{\pm.01}$ $0.18_{\pm.01}$
DiffS.-lb $0.19_{\pm.00}$ $0.43_{\pm.01}$ $0.56_{\pm.01}$ $0.43_{\pm.01}$ $0.54_{\pm.01}$ $0.73_{\pm.01}$ $48_{\pm.01} 0.0$$09_{\pm.00}$ $0.19_{\pm.01}$ $0.29_{\pm.01}$ $0.76_{\pm.01}$ $0.82_{\pm.00}$ $0.92_{\pm.00}$ $0.75_{\pm.01}$
DiffS.-minp $0.21_{\pm.00}$ $0.46_{\pm.01}$ $0.59_{\pm.01}$ $0.37_{\pm.01}$ $0.47_{\pm.01}$ $0.69_{\pm.01}$ $0.42_{\pm.01}$ $0.19_{\pm.00}$ $0.42_{\pm.01}$ $0.58_{\pm.01}$ $0.72_{\pm.01}$ $0.53_{\pm.00}$ $0.72_{\pm.01}$ $0.45_{\pm.01}$
@@
Top-$p$ $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.74_{\pm.01}$ $0.64_{\pm.00}$ $0.93_{\pm.00}$ $92_{\pm.00} 0.0$$0_{\pm.00} 0.0$$03_{\pm.00}$ $0.13_{\pm.00}$ $0.78_{\pm.01}$ $0.61_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$
$\eta$-sampling $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.11_{\pm.00}$ $0.74_{\pm.01}$ $0.64_{\pm.00}$ $0.93_{\pm.00}$ $93_{\pm.00} 0.0$$0_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.63_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$
Locally typical $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.63_{\pm.00}$ $0.93_{\pm.00}$ $93_{\pm.00} 0.0$$0_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.77_{\pm.01}$ $0.61_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$
Min-$p$ $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.11_{\pm.00}$ $0.74_{\pm.01}$ $0.64_{\pm.00}$ $0.93_{\pm.00}$ $93_{\pm.00} 0.0$$0_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.63_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$
DiffS.-cut $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.47_{\pm.01}$ $0.25_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.01}$ $0.39_{\pm.01}$ $0.19_{\pm.00}$ $0.34_{\pm.01}$ $0.19_{\pm.01}$
DiffS.-lb $0.17_{\pm.00}$ $0.40_{\pm.01}$ $0.52_{\pm.01}$ $0.48_{\pm.01}$ $0.59_{\pm.01}$ $0.78_{\pm.01}$ $54_{\pm.01} 0.0$$05_{\pm.00}$ $0.13_{\pm.00}$ $0.23_{\pm.01}$ $0.77_{\pm.01}$ $0.83_{\pm.00}$ $0.95_{\pm.00}$ $0.83_{\pm.01}$
DiffS.-minp $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.58_{\pm.01}$ $0.38_{\pm.01}$ $0.49_{\pm.01}$ $0.72_{\pm.01}$ $0.45_{\pm.01}$ $0.18_{\pm.00}$ $0.40_{\pm.01}$ $0.55_{\pm.01}$ $0.76_{\pm.01}$ $0.57_{\pm.00}$ $0.79_{\pm.01}$ $0.50_{\pm.01}$